# ArrowGEV: Grounding Events in Video via Learning the Arrow of Time

### v1

@@ Abstract
Grounding events in videos serves as a fundamental capability in video analysis. While Vision-Language Models (VLMs) are increasingly employed for this task, existing approaches predominantly train models to associate events with timestamps in the forward video only. This paradigm hinders VLMs from capturing the inherent temporal structure and directionality of events, thereby limiting robustness and generalization. To address this limitation, inspired by the arrow of time in physics, which characterizes the intrinsic directionality of temporal processes, we propose ArrowGEV, a reinforcement learning framework that explicitly models temporal directionality in events to improve both event grounding and temporal directionality understanding in VLMs. Specifically, we categorize events into time-sensitive (e.g., putting down a bag) and time-insensitive (e.g., holding a towel in the left hand). The former denote events whose reversal substantially alters their meaning, while the latter remain semantically unchanged under reversal. For time-sensitive events, ArrowGEV introduces a reward that encourages VLMs to discriminate between forward and backward videos, whereas for time-insensitive events, it enforces consistent grounding across both directions. Extensive experiments demonstrate that ArrowGEV not only improves grounding precision and temporal directionality recognition, but also enhances general video understanding and reasoning ability.
@@
[Uncaptioned image]Code   [Uncaptioned image]Model   [Uncaptioned image]Data Fangxu Yu1thanks: This work was done during the internship at Pattern Recognition Center, WeChat AI, Tencent Inc, China., Ziyao Lu2thanks: Corresponding author., Liqiang Niu2, Fandong Meng2, Jie Zhou2 1School of Artificial Intelligence, Nanjing University, China 2Pattern Recognition Center, WeChat AI, Tencent Inc, China yufx@smail.nju.edu.cn  fandongmeng@tencent.com
Figure 1: Two examples of Qwen2.5-VL-7B predicting event timestamps in forward and backward videos. In the top row, reversing the video changes the event semantics, while in the bottom row the event remains invariant. The model partially localizes events in forward videos but fails to recognize event absence in the first reversed case and cannot robustly localize the event in the second reversed video.
@@ 1 Introduction
To tackle this challenge, early approaches relied on handcrafted architectures and video-query feature-matching strategies (Hou et al., 2022; Pan et al., 2023), but they suffer from constraints of predefined video snippets, suboptimal video-text features, and poor cross-task generalization. Recent work has shifted to end-to-end Vision Language Models (VLMs) (Wang et al., 2025a; Team et al., 2025; Xiaomi, 2025), which directly process videos and queries while retaining generalizability either via large-scale timestamp annotation training (Huang et al., 2024a; Li et al., 2024d; Zeng et al., 2025), integration of textual timestamp tokens/embeddings (Guo et al., 2025), or adaptation to event grounding via video segmentation (Guo et al., 2024b; Wang et al., 2024). Despite this progress, existing methods only align events with forward videos, failing to capture the intrinsic temporal structure and directionality of events.
@@
To address this, we turn to the Arrow of Time (Eddington, 2019; Layzer, 1975), a foundational concept in physics that characterizes the intrinsic directionality of temporal processes. This perspective highlights that the semantics of real-world events are inherently tied to their temporal progression: reversing time may either yield a semantically distinct event or preserve the core meaning, depending on the nature of the process. Generally, the impact of this temporal directionality on event understanding manifests in two distinct event types: time-sensitive events, where reversal fundamentally alters meaning (e.g., "a man picks up a glass" becomes "a man puts down a glass"), and time-insensitive events, whose meaning remains invariant under reversal (e.g., "a ball is on the table" retains the same semantic essence when time is reversed).
@@
We conduct extensive experiments on three GEV benchmark datasets, results indicate that VLMs trained with ArrowGEV significantly improve the event grounding performance. In addition, ArrowGEV Substantially improved the VLM’s ability to understand temporal structures in event grounding. Finally, ArrowGEV also improves the out-of-distribution (OOD) performance on general video understanding and reasoning tasks.
@@ 2 Related Work
Grounding Events in Videos with VLMs. This task localizes specific events within untrimmed videos (Nan et al., 2021; Wang et al., 2018; Li et al., 2020; Zhao et al., 2017; Kulkarni and Fazli, 2025; Chen et al., 2025b; Yang et al., 2025b; Tian et al., 2025; Hannan et al., 2024; Mu et al., 2024b). While traditional methods rely on task-specific heads, recent Vision-Language Models (VLMs) leverage the reasoning capabilities of LLMs for unified temporal understanding (Huang et al., 2024a; Li et al., 2024d). To bridge the modality gap, existing research typically focuses on large-scale supervised fine-tuning with timestamp-based data (Zeng et al., 2025), introducing specialized temporal tokens (Hong et al., 2024), or utilizing video segmentation to align structural granularity (Huang et al., 2024b; Wang et al., 2024). Despite these advances, current VLM-based approaches largely overlook the intrinsic temporal directionality of events. In contrast, ArrowGEV introduces a principled framework that explicitly models temporal direction, facilitating more robust and physically consistent grounding.
@@
Post-Training for VLMs. Post-training techniques are essential for adapting pre-trained VLMs to complex downstream tasks. While large-scale instruction tuning has significantly boosted performance in models like LLaVA-OV and MAmmoTH-VL (Li et al., 2024a; Guo et al., 2024a), recent research has pivoted toward RL to refine multimodal reasoning (Yu et al., 2024; Su et al., 2025; Meng et al., 2025). This shift is particularly evident in video understanding, where RL is increasingly employed to enhance event grounding and long-form reasoning (Feng et al., 2025; Wang et al., 2025b). However, existing RL approaches for videos often overlook the inherent temporal structure of events. Our work addresses this gap by explicitly instilling temporal directionality, guiding VLMs toward a more robust and physically grounded understanding of videos.
@@
Time in Video. Temporal directionality is a foundational self-supervised signal for video representation learning, typically utilized through shuffle-and-learn or order-prediction tasks (Misra et al., 2016; Wei et al., 2018; Dwibedi et al., 2019; Yu et al., 2025). VLM frameworks often treat reversed videos merely as negative samples for contrastive alignment or as binary classification targets (Xu et al., 2021; Price and Damen, 2019). However, recent evidence suggests that VLMs remain alarmingly insensitive to temporal directionality in complex reasoning tasks (Xue et al., 2025; Du et al., 2024). Unlike prior work that focuses on events whose semantics change under reversal, we study temporally invariant events and move beyond coarse classification. We examine temporal directionality under the high-precision demands of event grounding, requiring models not only to detect reversal but also to localize events along the temporal axis accurately.
@@
Figure 2: Quatitative Analysis of Qwen2.5-VL-7B on GEV Benchmarks. Left: statistics of the portion of time-sensitive and time-insensitive events across three benchmarks. Right: R1@m metrics on the time-sensitive subsets of three benchmarks.
@@ 3.1 Pilot Study
To assess the ability of current Vision-Language Models (VLMs) to perceive temporal structure, we conduct a pilot study using the Qwen2.5-7B-VL-Instruct model. Our investigation focuses on time-sensitive events whose semantic meaning is fundamentally altered upon time reversal (e.g., "opening a door" becomes "closing a door"). As illustrated in Figure 2, such events constitute a significant portion of common benchmarks, particularly Charades-STA (Sigurdsson et al., 2016).
We evaluate the VLM by prompting it to localize time-sensitive events in both forward and backward videos. We then measure localization accuracy using Intersection over Union (IoU). For the forward video, we use the original ground truth, while for the backward video, we use a corresponding pseudo-ground truth created by reversing the original event’s timestamps. Ideally, a model with a robust grasp of temporal structure and directionality should recognize that the described event no longer exists in the reversed sequence, resulting in an Intersection-over-Union (IoU) score close to zero. However, as shown in the right panel of Figure 2, the VLM localizes these time-sensitive events even in backward videos, producing a high IoU. This reveals a critical failure: the model fails to capture temporal structure, struggling to distinguish the semantic change of events in forward and backward video.
@@ 3.2 Background of GRPO: RL for LLM
As a pioneer among open-sourced R1-style LLMs, Deepseek-R1 (9) leverages Group Relative Policy Optimization (GRPO) to train the policy model $\pi_{\theta}$ (i.e., the LLM) to think before answering, making it particularly well-suited for tasks with well-defined answers, such as mathematical reasoning. In the GRPO framework, given an input question $p$, the LLM samples $G$ candidate responses $o=\{o_{1},\dots,o_{G}\}$, and a reward function $r(\cdot)$ assigns a reward score to each response, yielding $\{r(o_{1}),\dots,r(o_{G})\}$. GRPO encourages the LLM to generate responses that maximize a weighted sum reward $R(o)$, defined by:
@@
where $\pi_{\theta}(o)$ denotes the probability of LLM generating the response $o$, and $\pi_{\theta_{\mathrm{old}}}$ represents the LLM parameters from a recently optimized state. And the latter term is the Advantage $A_{i}$ of $i$-th candidate. To ensure training stability and avoid large deviations from the original language model behavior, the final training objective incorporates a KL-divergence regularization term (9), penalizing divergence between $\pi_{\theta}$ and $\pi_{\mathrm{old}}$:
@@ 3.3 ArrowGEV
Formulation. The task of Grounding events in videos requires a model $\mathcal{M}$ to map an untrimmed video $V$ and a natural language event $\mathcal{q}$ to a specific temporal timestamp $\mathcal{T}^{\text{fwd}}=\mathcal{M}(\mathcal{V},\mathcal{q})$ that accurately aligns with a ground-truth annotation $\mathcal{T}^{\text{gt}}$. As discussed in Section 3.1, the VLM struggles to capture the temporal structure of events in the video. To overcome this limitation, we leverage the backward video, denoted as $\mathcal{V}^{\prime}$, to enhance the understanding of temporal structure. Our key insight is that the effect of temporal reversal is not uniform across all events; it depends on the event’s intrinsic properties. Therefore, to construct a meaningful learning signal from the reversed video $V^{\prime}$, we first categorize events according to their temporal nature.
@@
Then, we design an RL framework to encourage the VLM to not only accurately localize the timestamps of the events in the forward video, but also to distinguish the existence of time-sensitive events and recognize time-insensitive events in the backward video. After training, ArrowGEV enables the VLM achieve above and learn the temporal structure more robustly to enhance the localization.
@@ 3.4 Temporal Directionality Reward Modeling
A reward function for GEV should ideally incentivize a model to localize the events while understanding an event’s temporal structure.
To determine the type of each event, we use an LLM to perform reasoning to classify each event $q$ as time-sensitive or time-insensitive. The model first generates the reasoning process $r\sim P_{\text{LLM}}(\cdot|p,q)$, and then determines the type of event $c(q)\sim P_{\text{LLM}}(\cdot|r,p,q)$. The analysis of event categorization verifies the reliability in the Appendix A, and see the prompt in the Appendix E.
@@
The accuracy reward $r_{\text{acc}}$ is defined as timestamp-aware IoU (Wang et al., 2025c) with the ground-truth $\mathcal{T}^{\text{gt}}$, which encourages the alignment between start and end time with the ground-truth based on the standard IoU:
$\mathrm{tIoU}(\mathcal{T}^{\text{fwd}},\mathcal{T}^{\text{gt}})=\text{IoU}\cdot(1-\frac{|t_{s}-t^{\prime}_{s}|}{t})\cdot(1-\frac{|t_{e}-t^{\prime}_{e}|}{t})$ (4)
@@
To define the temporal directionality reward, let $\mathcal{T}^{\text{rev}}=\mathcal{M}(V^{\prime},q)$ be the VLM’s predicted timestamp on the backward video. We define a temporal reversal operator $\mathcal{R}(\mathcal{T})\triangleq[d-t_{e},d-t_{s}]$ for a timestamp $\mathcal{T}=[t_{s},t_{e}]$ in a video of duration $d$. This allows us to compute a directionality score, $S_{c}$, measuring the alignment between the reversed prediction and the mirrored forward prediction:
@@ 3.5 Training
Difficulty-Aware Weight Adjustment. During the training, samples progressively become easier for the model. This will lead to a difficulty-bias issue. To alleviate this, we propose a weight adjustment that enables the model to focus more on the hard samples. Specifically, we introduce a difficulty coefficient $w_{i}\propto-\text{tIoU}(\mathcal{T}^{\text{fwd}},\mathcal{T}^{\text{gt}})$ for $i$-th sample to quantify the difficulty level. In this case, we sample $G$ predictions by the VLM and calculate the difficulty weight:
@@
The efficacy of the above detailed policy optimization is highly dependent on the quality and challenge of the training data distribution, $\mathcal{D}$. As the policy VLM $\pi_{\theta}$ improves, a static data set can be dominated by samples that no longer provide a sufficient learning signal. To counteract this and maintain a challenging training environment, we implement a dynamic curriculum strategy that adaptively refines the data distribution throughout the training. Specifically, we initialize our training set, $\mathcal{D}_{0}$, with the dataset from (Wang et al., 2025c), which is pre-filtered to emphasize samples of moderate difficulty. To ensure the model is persistently challenged as it learns, we introduce a dynamic difficulty filter at the conclusion of each training epoch $e$. To construct the dataset for the next epoch, $\mathcal{D}_{e+1}$, we evaluate each sample $(\mathcal{V},\mathcal{q})\in\mathcal{D}_{e}$ against the current policy $\pi_{\theta_{e}}$. A sample is deemed "mastered" and is subsequently removed if the policy consistently solves it with high accuracy. Formally, for each $(\mathcal{V},\mathcal{q})$, we generate a group of $G$ rollout outputs $\{\mathcal{T}_{i}^{fwd}\}_{i=1}^{G}$. The sample is filtered out if its worst-case performance in the group exceeds a high-performance threshold $\eta$:
$\mathcal{D}_{e+1}=\mathcal{D}_{e}\setminus\left\{(\mathcal{V},\mathcal{q})\in\mathcal{D}_{e}\mid\min_{i=1\dots G}\mathrm{IoU}(\mathcal{T}_{i}^{fwd},\mathcal{T}^{\text{gt}})>\eta\right\}$ (10)
We set $\eta=0.7$ in our experiments, which is the most strict metric for evaluating the quality of prediction in most of the previous work (Ye et al., 2025; Li et al., 2025; Nguyen et al., 2025). This adaptive curriculum ensures that the model continually focuses its capacity on unsolved or challenging problems, thereby maintaining a strong gradient signal and promoting the development of a more robust and generalizable policy.

### v2

@@ Abstract
Grounding events in videos serves as a fundamental capability in video analysis. While Vision Language Models (VLMs) are increasingly employed for this task, existing approaches predominantly train models to associate events with timestamps in the forward video only. This paradigm hinders VLMs from capturing the inherent temporal structure and directionality of events, thereby limiting robustness and generalization. To address this limitation, inspired by the arrow of time in physics, which characterizes the intrinsic directionality of temporal processes, we propose ArrowGEV, a reinforcement learning framework that explicitly models temporal directionality in events to improve both event grounding and temporal directionality understanding in VLMs. Specifically, we categorize events into time-sensitive (e.g., putting down a bag) and time-insensitive (e.g., holding a towel in the left hand). The former denote events whose reversal substantially alters their meaning, while the latter remain semantically unchanged under reversal. For time-sensitive events, ArrowGEV introduces a reward that encourages VLMs to discriminate between forward and backward videos, whereas for time-insensitive events, it enforces consistent grounding across both directions. Extensive experiments demonstrate that ArrowGEV not only improves grounding precision and temporal directionality recognition, but also enhances general video understanding and reasoning ability.
@@
[Uncaptioned image]Code   [Uncaptioned image]Model   [Uncaptioned image]Data Fangxu Yu1thanks: This work was done during the internship at Pattern Recognition Center, WeChat AI, Tencent Inc, China., Ziyao Lu2thanks: Corresponding author., Liqiang Niu2, Fandong Meng2, Jie Zhou2 1School of Artificial Intelligence, Nanjing University, China 2Pattern Recognition Center, WeChat AI, Tencent Inc, China yufx@smail.nju.edu.cn  ziyaolu@tencent.com
Figure 1: Two examples of Qwen2.5-VL-7B predicting event timestamps in forward and backward videos. In the top row, reversing the video changes the event semantics, while in the bottom row, the event remains invariant. The model partially localizes events in forward videos but fails to recognize event absence in the first reversed case and cannot robustly localize the event in the second backward video.
@@ 1 Introduction
To tackle this challenge, early approaches relied on handcrafted architectures and video-query feature-matching strategies (Hou et al., 2022; Pan et al., 2023), but they suffer from constraints of predefined video snippets, suboptimal video-text features, and poor cross-task generalization. Recent work has shifted to end-to-end Vision Language Models (VLMs) (Wang et al., 2025a; Team et al., 2025; Xiaomi, 2025), which directly process videos and queries while retaining generalizability either via large-scale timestamp annotation training (Huang et al., 2024a; Li et al., 2024d; Zeng et al., 2025), integration of textual timestamp tokens/embeddings (Guo et al., 2025b), or adaptation to event grounding via video segmentation (Guo et al., 2024b; Wang et al., 2024). Despite this progress, existing methods only align events with forward videos, failing to capture the intrinsic temporal structure and directionality of events.
@@
To address this, we turn to the Arrow of Time (Eddington, 2019; Layzer, 1975), a foundational concept in physics characterizing the intrinsic directionality of temporal processes. This perspective highlights that the semantics of real-world events are inherently tied to their temporal progression: reversing time may either yield a semantically distinct event or preserve the core meaning, depending on the nature of the process. The impact of this temporal directionality on event understanding manifests itself in two distinct event types: time-sensitive events, where reversal fundamentally alters meaning (e.g., "a man picks up a glass" becomes "a man puts down a glass"), and time-insensitive events, whose meaning remains invariant under reversal (e.g., "a ball is on the table" retains the same semantics when time is reversed).
@@
We conduct extensive experiments on three GEV benchmark datasets, and results indicate that VLMs trained with ArrowGEV significantly improve the event grounding performance. In addition, ArrowGEV substantially improved the VLM’s ability to understand temporal structures in event grounding. Finally, ArrowGEV also improves the out-of-distribution (OOD) performance on general video understanding and reasoning tasks.
@@ 2 Related Work
Grounding Events in Videos with VLMs. This task localizes specific events within untrimmed videos (Nan et al., 2021; Wang et al., 2018; Li et al., 2020; Zhao et al., 2017; Kulkarni and Fazli, 2025; Chen et al., 2025c; Yang et al., 2025b; Tian et al., 2025; Hannan et al., 2024; Mu et al., 2024b; Kim et al., 2024; Oncescu et al., 2021). Although traditional methods rely on task-specific heads, recent Vision Language Models (VLMs) leverage the reasoning capabilities of LLMs for unified temporal understanding (Huang et al., 2024a; Li et al., 2024d). To bridge the modality gap, existing research typically focuses on large-scale supervised fine-tuning with timestamp-based data (Zeng et al., 2025), introducing specialized temporal tokens (Hong et al., 2024), or utilizing video segmentation to align structural granularity (Huang et al., 2024b; Wang et al., 2024). Despite these advances, current VLM-based approaches largely overlook the intrinsic temporal directionality of events. In contrast, ArrowGEV introduces a principled framework that explicitly models temporal direction, facilitating more robust and physically consistent grounding.
@@
Post-Training for VLMs. Post-training techniques are essential for adapting pre-trained VLMs to complex downstream tasks. Although large-scale instruction tuning has significantly improved performance in models such as LLaVA-OV (Li et al., 2024a) and MAmmoTH-VL (Guo et al., 2024a), recent research has been geared toward RL to refine multimodal reasoning (Yu et al., 2024; Su et al., 2025; Meng et al., 2025; Chen et al., 2025a). This shift is particularly evident in video understanding, where RL is increasingly employed to enhance event grounding and long-form reasoning (Feng et al., 2025; Wang et al., 2025b). However, existing RL approaches for videos often overlook the inherent temporal structure of events. Our work addresses this gap by explicitly instilling temporal directionality, guiding VLMs toward a more robust and physically grounded understanding of videos.
@@
Time in Video. Temporal directionality is a foundational self-supervised signal for video representation learning, typically utilized through shuffle-and-learn or order-prediction tasks (Misra et al., 2016; Wei et al., 2018; Dwibedi et al., 2019; Yu et al., 2025). VLM frameworks often treat backward videos only as negative samples for contrastive alignment or as binary classification targets (Xu et al., 2021; Price and Damen, 2019). However, recent evidence suggests that VLMs remain insensitive to temporal directionality in complex reasoning tasks (Xue et al., 2025; Du et al., 2024). Unlike prior work that focuses on events whose semantics change under reversal, we study both time-sensitive and time-invariant events and move beyond coarse classification. We examine temporal directionality under the high-precision demands of event grounding, requiring models not only to detect reversal but also to localize events along the temporal axis.
@@
Figure 2: Quantitative Analysis of Qwen2.5-VL-7B on GEV Benchmarks. Left: statistics of the portion of time-sensitive and time-insensitive events across three benchmarks. Right: R1@m metrics on the time-sensitive subsets of three benchmarks.
@@ 3.1 Pilot Study
To assess the ability of current Vision Language Models (VLMs) to perceive temporal structure, we conduct a pilot study using the Qwen2.5-7B-VL-Instruct model. Our investigation focuses on time-sensitive events whose semantic meaning is fundamentally altered upon time reversal (e.g., "opening a door" becomes "closing a door"). As illustrated in Figure 2, such events constitute a significant portion of common benchmarks, particularly Charades-STA (Sigurdsson et al., 2016).
We evaluate the VLM by prompting it to localize time-sensitive events in both forward and backward videos. We then measure localization accuracy using Intersection over Union (IoU). For the forward video, we use the original ground truth, while for the backward video, we use a corresponding pseudo-ground truth created by reversing the original event’s timestamps. Ideally, a model with a robust grasp of temporal structure and directionality should recognize that the described event no longer exists in the reversed sequence, resulting in an IoU score close to zero. However, as shown in the right panel of Figure 2, the VLM localizes these time-sensitive events even in backward videos, producing a non-trivial IoU. This reveals a critical failure: the model fails to capture temporal structure, struggling to distinguish the semantic change of events in forward and backward video.
@@ 3.2 Background of GRPO: RL for LLM
As a pioneer among open-source R1-style LLMs, DeepSeek-R1 (Guo et al., 2025a) leverages Group Relative Policy Optimization (GRPO) to train the policy model $\pi_{\theta}$ (i.e., the LLM) to think before answering, making it particularly suitable for tasks with well-defined answers, such as mathematical reasoning. In the GRPO framework, given an input question $p$, the LLM samples $G$ candidate responses $o=\{o_{1},\dots,o_{G}\}$, and a reward function $r(\cdot)$ assigns a reward score to each response, yielding $\{r(o_{1}),\dots,r(o_{G})\}$. GRPO encourages the LLM to generate responses that maximize a weighted sum reward $R(o)$, defined by:
@@
where $\pi_{\theta}(o)$ denotes the probability of LLM generating the response $o$, and $\pi_{\theta_{\mathrm{old}}}$ represents the LLM parameters from a recently optimized state. And the latter term is the Advantage $A_{i}$ of $i$-th candidate. To ensure training stability and avoid large deviations from the original language model behavior, the final training objective incorporates a KL-divergence regularization term (Guo et al., 2025a), penalizing divergence between $\pi_{\theta}$ and $\pi_{\mathrm{old}}$:
@@ 3.3 ArrowGEV
Formulation. The task of grounding events in videos requires a model $\mathcal{M}$ to map an untrimmed video $\mathcal{V}$ of duration $d$ and a natural language event $q$ to a specific temporal timestamp $\mathcal{T}^{\text{fwd}}=\mathcal{M}(\mathcal{V},q)$ that accurately aligns with a ground-truth annotation $\mathcal{T}^{\text{gt}}$. As discussed in Section 3.1, the VLM struggles to capture the temporal structure of events in the video. To overcome this limitation, we leverage the backward video, denoted as $\mathcal{V}^{\prime}$, to enhance the understanding of temporal structure. Our key insight is that the effect of temporal reversal is not uniform across all events; it depends on the event’s intrinsic properties. Therefore, to construct a meaningful learning signal from the backward video $\mathcal{V}^{\prime}$, we first categorize events according to their temporal nature.
@@
Then, we design an RL framework to encourage the VLM to not only accurately localize the timestamps of the events in the forward video, but also to distinguish the existence of time-sensitive events and recognize time-insensitive events in the backward video. After training, ArrowGEV enables the VLM to achieve above and learn the temporal structure more robustly to enhance the localization.
@@ 3.4 Temporal Directionality Reward Modeling
A reward function for GEV should ideally incentivize a model to localize events while understanding an event’s temporal structure.
To determine the type of each event, we use an LLM to perform reasoning to classify each event $q$ as time-sensitive or time-insensitive with prompt $p$. The model first generates the reasoning process $r\sim P_{\text{LLM}}(\cdot|p,q)$, and then determines the type of event $c(q)\sim P_{\text{LLM}}(\cdot|r,p,q)$. The analysis in Appendix A verifies the reliability of event categorization. See the prompt in Appendix E.
@@
The accuracy reward $r_{\text{acc}}$ is defined as timestamp-aware IoU (Wang et al., 2025b) with the ground-truth $\mathcal{T}^{\text{gt}}$, which encourages the alignment between start and end time with the ground-truth based on the standard IoU:
$\mathrm{tIoU}(\mathcal{T}^{\text{fwd}},\mathcal{T}^{\text{gt}})=\text{IoU}\cdot(1-\frac{|t_{s}-t^{\prime}_{s}|}{d})\cdot(1-\frac{|t_{e}-t^{\prime}_{e}|}{d})$ (4)
@@
To define the temporal directionality reward, let $\mathcal{T}^{\text{rev}}=\mathcal{M}(\mathcal{V}^{\prime},q)$ be the VLM’s predicted timestamp on the backward video. We define a temporal reversal operator $\mathcal{R}(\mathcal{T})\triangleq[d-t_{e},d-t_{s}]$ for a timestamp $\mathcal{T}=[t_{s},t_{e}]$ in a video. This allows us to compute a directionality score, $S_{c}$, measuring the alignment between the reversed prediction and the mirrored forward prediction:
@@ 3.5 Training
Difficulty-Aware Weight Adjustment. During training, samples become progressively easier for the model. This will lead to a difficulty-bias issue. To alleviate this, we propose a weight adjustment that enables the model to focus more on the hard samples. Specifically, we introduce a difficulty coefficient $w_{i}\propto-\text{tIoU}(\mathcal{T}^{\text{fwd}},\mathcal{T}^{\text{gt}})$ for $i$-th sample to quantify the difficulty level. In this case, we sample $G$ predictions by the VLM and calculate the difficulty weight:
@@
The efficacy of the above-detailed policy optimization is highly dependent on the quality and challenge of the training data distribution, $\mathcal{D}$. As the policy VLM $\pi_{\theta}$ improves, a static dataset can be dominated by samples that no longer provide a sufficient learning signal. To counteract this and maintain a challenging training environment, we implement a dynamic curriculum strategy that adaptively refines the data distribution throughout the training. Specifically, we initialize our training set, $\mathcal{D}_{0}$, with the dataset from (Wang et al., 2025b), which is pre-filtered to emphasize samples of moderate difficulty. To ensure the model is persistently challenged as it learns, we introduce a dynamic difficulty filter at the conclusion of each training epoch $e$. To construct the dataset for the next epoch, $\mathcal{D}_{e+1}$, we evaluate each sample $(\mathcal{V},q)\in\mathcal{D}_{e}$ against the current policy $\pi_{\theta_{e}}$. A sample is deemed "mastered" and is subsequently removed if the policy consistently solves it with high accuracy. Formally, for each $(\mathcal{V},q)$, we generate a group of $G$ rollout outputs $\{\mathcal{T}_{i}^{fwd}\}_{i=1}^{G}$. The sample is filtered out if its worst-case performance in the group exceeds a high-performance threshold $\eta$:
$\mathcal{D}_{e+1}=\mathcal{D}_{e}\setminus\left\{(\mathcal{V},q)\in\mathcal{D}_{e}\mid\min_{i=1\dots G}\mathrm{IoU}(\mathcal{T}_{i}^{fwd},\mathcal{T}^{\text{gt}})>\eta\right\}$ (10)
We set $\eta=0.7$ in our experiments, which is the strictest IoU threshold commonly adopted in prior work (Wang et al., 2025b; Li et al., 2025; Nguyen et al., 2025). This adaptive curriculum ensures that the model continually focuses its capacity on unsolved or challenging problems, thereby maintaining a strong gradient signal and promoting the development of a more robust and generalizable policy.